OpenStudy (anonymous):

Solve for x. log(7)*x+log(7)*(x-1)=log(7)*2x

OpenStudy (anonymous):

you'll get this :\[\log_{7}x(x-1) = \log_{7} 2x \] but I'm not sure whether you're supposed to take the ln of both sides to get rid of the log

OpenStudy (anonymous):

when we have (+) and same log bases, then it means multiplication, and when we have (-) and same log bases then we divide ^_^

OpenStudy (anonymous):

but if you do take the ln of both sides then you'll get :\[x^2-x=2x \rightarrow x^2-3x=0\]\[x(x-3) = 0\] so x = 0, 3 ^_^ but I'm not sure though

OpenStudy (anonymous):

wow i was totally over thinking it lol thank you sooo much

OpenStudy (anonymous):

np ^_^, but again not sure with the ln part :)